OpenStudy (anonymous):

Hello, do you know where I can find the standard deviation of the S&P 500 index for 2010

OpenStudy (amistre64):

Standard deviation can be found an i number of different ways: One being: \[\sigma=\sqrt{\frac{\color{green}{\sum}{\color{purple}{(x-\mu)^2}}}{\color{red}{N}}}\] for large amounts of data tho a more convenient form is this: \[s=\sqrt{\frac{n(\sum{(x^2)})-(\sum{x})^2}{n(n-1)}}\] where n=the number of items in your data. and x=each value of data

OpenStudy (amistre64):

where to find this already? dunno, google maybe

OpenStudy (anonymous):

http://www.moneychimp.com/features/market_cagr.htm